# Effective bias expansion for circumventing 21 cm foregrounds

### v1

@@ I Introduction
The time from the birth of the first stars to the end of the Epoch of Reionization (EoR) marks an important era for the astrophysics of galaxy formation and assembly, as well as for the underlying growth of cosmological structure Loeb and Barkana (2001); Zaroubi (2013); Robertson (2022); McQuinn et al. (2007); Mesinger (2016); Liu et al. (2016). The redshifted 21 cm line from neutral hydrogen is a direct probe of this period of cosmic time Furlanetto et al. (2006); Pritchard and Loeb (2012), which is being actively targeted by many existing and proposed radio observatories Monsalve et al. (2017); Nambissan T. et al. (2021); Philip et al. (2019); Parsons et al. (2010); van Haarlem et al. (2013); Zarka et al. (2012); Tingay et al. (2013); Wayth et al. (2018); Gupta et al. (2017); DeBoer et al. (2017); Berkhout et al. (2024); Eastwood et al. (2019); Koopmans et al. (2015). In particular, radio interferometers have already set stringent upper bounds on the 21 cm power spectrum (Parsons et al., 2014; Dillon et al., 2014, 2015; Ewall-Wice et al., 2016; Beardsley et al., 2016; Patil et al., 2017; Li et al., 2019; Barry et al., 2019; Eastwood et al., 2019; Gehlot et al., 2019; Kolopanis et al., 2019; Trott et al., 2020; Mertens et al., 2020; Garsden et al., 2021; Yoshiura et al., 2021; Abdurashidova et al., 2022a; HERA Collaboration et al., 2023; Wilensky et al., 2023; Munshi et al., 2024; Acharya et al., 2024; Mertens et al., 2025; Nunhokee et al., 2025), which have been used to constrain the properties of the intergalactic medium (IGM) at high redshifts Abdurashidova et al. (2022b); HERA Collaboration et al. (2023); Ghara et al. (2025); Nunhokee et al. (2025).
@@
A major ongoing issue in 21 cm cosmology is that the signal is highly contaminated by foreground emission that is many orders of magnitude brighter (see Ref. Liu and Shaw (2020) for a review). These foregrounds, which include Galactic synchrotron radiation, are expected to be spectrally smooth and will therefore affect large-scale modes of the 21 cm signal along the line of sight. Additionally, the chromatic response of interferometers can cause the foregrounds to contaminate other Fourier modes. Decomposing each wavevector into $\mathbf{k}\equiv(\mathbf{k}_{\perp},k_{\parallel})$, where $k_{\parallel}$ denotes the Fourier modes parallel to the line of sight and $\mathbf{k}_{\perp}$ represents the Fourier modes in the plane perpendicular to the line of sight, then in $_{\perp}-kk_{\parallel}$ space, the region where foreground contamination occurs is referred to as the “foreground wedge” due to its shape Datta et al. (2010); Morales et al. (2012); Parsons et al. (2012); Thyagarajan et al. (2015); Liu et al. (2014a, b); Munshi et al. (2025).
@@
It is difficult to completely subtract foregrounds Datta et al. (2010); Liu and Tegmark (2011); Chapman et al. (2012); Ghosh et al. (2015); Ewall-Wice et al. (2021); Mertens et al. (2018); Wang et al. (2024) in the wedge due to uncertainties in modeling both the foreground emission and the chromatic response of the instrument Barry et al. (2016); Patil et al. (2016); Ewall-Wice et al. (2017); Byrne et al. (2019); Mouri Sardarabadi and Koopmans (2019); Neben et al. (2016); Joseph et al. (2018); Ansah-Narh et al. (2018); Zheng et al. (2017); Orosz et al. (2019); Joseph et al. (2020); Choudhuri et al. (2021); Kim et al. (2022); Ewall-Wice et al. (2016); Kern et al. (2019); Ung et al. (2020); Josaitis et al. (2022); Rath et al. (2024); Chen et al. (2025). Therefore, these modes are often discarded, with many analyses of 21 cm data focusing on evading foreground-contaminated modes Vedantham et al. (2012); Trott et al. (2012); Hazelton et al. (2013); Thyagarajan et al. (2013); Liu et al. (2014). This effectively renders a large number of cosmological modes inaccessible to current experiments, significantly reducing their sensitivity. It also limits the possibility of correlating the 21 cm signal with complementary tracers of the matter field from high-redshift galaxy surveys or line intensity mapping experiments La Plante et al. (2023); Gagnon-Hartman et al. (2025); Fronenberg and Liu (2024).
@@
The rest of this paper is organized as follows. In Section II, we provide a conceptual overview of how we reconstruct missing modes and describe the various assumptions we make about what density information is available. In Section III, we briefly review the effective 21 cm bias expansion of Refs. McQuinn and D’Aloisio (2018); Qin et al. (2022) and describe our methods for reconstructing missing density modes. In Section V, we present the results of applying the bias expansion to different combinations of density information, and also demonstrate the complementarity of measurements from galaxy surveys. We conclude in Section VI. Throughout this article, we use cosmological parameters from Planck 2018 Aghanim et al. (2020).
@@ II Conceptual Overview
Devising an optimal strategy for extracting density information from 21 cm fluctuations would be worth its own dedicated study, and therefore we choose to focus primarily on steps 3 and 4 in this work. However, we note that Ref. Chen et al. (in preparation) has shown that step 2 is possible by adopting a gradient-based sampler to simultaneously infer the bias parameters and the underlying density field from the observed 21 cm modes outside the foreground wedge. Specifically, deep in the perturbative regime, the 21 cm field and density field are primarily related through the linear bias term which is local Qin et al. (2022, 2025). This indicates that the density modes that can be recovered correspond to the same modes where one has uncontaminated 21 cm information. This is corroborated by the findings of Ref. Chen et al. (in preparation), which showed explicitly that density modes outside the foreground wedge can be reliably inferred with up to $90\%$ accuracy on large scales (depending on the region of Fourier space) assuming a $\sim 3\sigma$ detection of those same individual 21 cm modes. As it turns out, the same approach of Ref. Chen et al. (in preparation) can also be used to reconstruct some of the large-scale density modes in the foreground wedge because their forward model includes nonlinear mode-coupling. These results and future improvements to step 2 can interface with the framework presented here in a straightforward way.
@@
  • the Fourier modes $\mathbf{k}$ outside of the foreground wedge, which are those with Liu et al. (2014)
@@
  • the modes $k_{\parallel}>k_{\mathrm{cut}}$, corresponding to a scenario where densely spaced or specifically arranged interferometer array configurations make it possible to use foreground mitigation strategies that may be able to recover wedge modes at high $k_{\perp}$ Murray and Trott (2018); MacKay et al. (in preparation). We also include this case since in our discussion because it enables a more direct comparison with results from the literature Darwish et al. (2021). We will refer to this as the “flat” case.
@@ IV Setup and mock fields
To evaluate the expression in Eq. (18), we first filter the nonlinear small scales from the tracer field. As above, we denote the smoothed tracer field by $\delta_{\mathrm{NL}}^{W}$. We calculate the wavenumber cutoff $k_{\mathrm{NL}}$ for this filter by smoothing $\delta_{\mathrm{NL}}$ with a Gaussian window function on increasingly large length scales until the largest overdensity is less than 0.8; the value of $k_{\mathrm{NL}}$ is not particularly sensitive to the choice of value that we use for largest overdensity or the choice of window function Qin et al. (2022). For the 21cmFAST realization used in this work, we find that the nonlinear scale is approximately $k_{\mathrm{NL}}=0.4$ cMpc-1 at z=8.5. The third column of Table 1 shows the number of uncontaminated modes that are in the perturbative regime, i.e. with $k<k_{\mathrm{NL}}$. To evaluate the total power spectra in Eq. (23), we use the CAMB Boltzmann code Lewis and Challinor (2011).
@@ V.1 Individual terms
Figure 5: The 21 cm signal resulting from substituting the density field with different modifications into Eq. 12. The first row shows the cross-correlation with the true 21 cm signal, $r_{\mathrm{21}}$, and the second row shows the ratio of the power spectra of the two fields. For the last three columns, we assume $\delta_{\mathrm{m,obs}}$ are the density modes lying outside of a wedge-shaped foreground region. In order to facilitate comparison of the different input density fields, we use the bias coefficients fit using the full density field across all columns.
@@ V.2 Adding density information from galaxy surveys
In Fig. 5, we show $r_{21}$ and $P_{\mathrm{total}}/P_{21}$ for the 21 cm signal calculated assuming wedge-shaped foregrounds using different assumptions about the underlying density field. In order to ensure that differences between each column are not merely due to changes in the best-fit bias parameters, in each panel we use the same bias parameters as determined from the full density field. The first column shows the result of substituting the full density field, $\delta_{\mathrm{m,full}}$, into Eq. (12). This is the best recovery of the 21 cm signal that we could ever obtain using the bias expansion. Here, we see that the cross-correlation is about 0.9 for wavenumbers within the perturbative regime. The amplitude of the power spectrum is also recovered with high accuracy and is within a factor of 1.5 of the true 21 cm power spectrum.
@@
The middle two columns serve as a comparison of the total reconstructed 21 cm field in the wedge with and without performing density reconstruction in the wedge. Hence, they are quite similar to the final columns of Figs. 2 and Figs. 3, but as explained above are constructed using a slightly different set of bias parameters. As discussed in Section V.1, the cross-correlation is slightly better without first performing density reconstruction. In other words, even if we assume that the matter density in the foreground wedge is zero, which is clearly incorrect, the mode-coupling intrinsic to the EFT terms is enough to yield a cross-correlation at the level of $r\sim 0.3$. This behavior can be understood as arising from the relative contribution of the linear terms with and without in-wedge density reconstruction, as described in Section V.1: the slight degradation of the cross-correlation is due to the fact that the linear terms in the bias expansion are less correlated with the true 21 cm fluctuations than the nonlinear terms. Additionally, without first performing density reconstruction, the reconstructed in-wedge 21 cm field underestimates the power spectrum by a factor of 8, due to the fact that we are missing contributions to the 21 cm fluctuations from many of the underlying density modes, whereas with density reconstruction the reconstructed 21 cm field overestimates the power spectrum by a factor of 90 (this is larger than what was reported in the previous section due to the different bias coefficients).
@@
The last column of Fig. 5 shows the result of substituting $\delta_{\mathrm{m,obs}}\oplus\delta_{\mathrm{m,Roman}}$ into Eq. (12). In this case, for modes that are within the foreground wedge and also not covered by the galaxy survey, the recovered cross-correlation is at the level of 0.5, and the power spectrum is within a factor of a few of the true amplitude. This is a clear improvement over using either $\delta_{\mathrm{m,obs}}$ or $\delta_{\mathrm{m,obs}}\oplus\delta_{\mathrm{m,recon}}$. This highlights the complementarity of high redshift galaxy surveys to 21cm experiments, and motivates their joint use in recovering the underlying density field as well as reconstruction of 21 cm modes obscured by foregrounds.
@@ VI Conclusion
Assuming that we have a method of obtaining density information outside of the foreground wedge Chen et al. (in preparation), we find that simply applying our bias expansion to these modes already yields $\sim 30\%$ cross-correlations with the true 21 cm signal within the foreground region, due to non-linear couplings between small- and large-scale modes. Applying the density reconstruction method of Ref. Darwish et al. (2021) to obtain additional density information within the foreground wedge improves the recovery of linear contributions to the 21 cm signal within the foreground region. For the wedge-shaped foreground scenario, density reconstruction does not improve the overall cross-correlation. Density reconstruction applied to the “flat” and “isotropic” scenarios yields results that are much improved compared to the density-reconstructed “wedge” case, giving cross-correlations as high as $70\%$ and power spectra that are within a factor of a few of the true value—this motivates further study of independent methods for extracting small-scale density information.
@@
In summary, we have demonstrated that bias expansions combined with EFT-based treatments of nonlinear mode-coupling provide a useful starting point for recovering 21 cm information lost to foregrounds. Their utility is enhanced when combined with reliable extractions of the underlying density field within the foreground region, e.g. from galaxy surveys. There are a number of potential directions for improvement on the results presented here. For example, the density estimator of Ref. Darwish et al. (2021) that we apply throughout is derived using an unrenormalized bias expansion; it would be instructive to rederive this estimator using renormalized operators and including all of the appropriate counterterms. Moreover, an assumption that we make throughout this work is that one can extract density information outside of the 21 cm foreground wedge from 21 cm measurements due to the strength of the linear bias term. However, this assumption merits a detailed study in its own right and new techniques to extract this density information, such as those presented in Ref. Chen et al. (in preparation), could be combined with the techniques studied here in a straightforward way. We leave these as directions for future study.

### v2

@@ I Introduction
The time from the birth of the first stars to the end of the Epoch of Reionization (EoR) marks an important era for the astrophysics of galaxy formation and assembly, as well as for the underlying growth of cosmological structure Loeb and Barkana (2001); Zaroubi (2013); Robertson (2022); McQuinn et al. (2007); Mesinger (2016); Liu et al. (2016). The redshifted 21 cm line from neutral hydrogen is a direct probe of this period of cosmic time Furlanetto et al. (2006); Pritchard and Loeb (2012), which is being actively targeted by many existing and proposed radio observatories Monsalve et al. (2017); Nambissan T. et al. (2021); Philip et al. (2019); Parsons et al. (2010); van Haarlem et al. (2013); Zarka et al. (2012); Tingay et al. (2013); Wayth et al. (2018); Gupta et al. (2017); DeBoer and others (2017); Berkhout et al. (2024); Eastwood et al. (2019); Koopmans et al. (2015). In particular, radio interferometers have already set stringent upper bounds on the 21 cm power spectrum (Parsons et al., 2014; Dillon et al., 2014, 2015; Ewall-Wice et al., 2016; Beardsley et al., 2016; Patil et al., 2017; Li et al., 2019; Barry et al., 2019; Eastwood et al., 2019; Gehlot et al., 2019; Kolopanis et al., 2019; Trott et al., 2020; Mertens and others, 2020; Garsden et al., 2021; Yoshiura et al., 2021; Abdurashidova et al., 2022b; HERA Collaboration et al., 2023; Wilensky et al., 2023; Munshi et al., 2024; Acharya et al., 2024; Mertens and others, 2025; Nunhokee and others, 2025; Abdurashidova and others, 2025), which have been used to constrain the properties of the intergalactic medium (IGM) at high redshifts Abdurashidova et al. (2022a); HERA Collaboration et al. (2023); Ghara and others (2025); Nunhokee and others (2025).
@@
A major ongoing issue in 21 cm cosmology is that the signal is highly contaminated by foreground emission that is many orders of magnitude brighter (see Ref. Liu and Shaw (2020) for a review). These foregrounds, which include Galactic synchrotron radiation, are expected to be spectrally smooth and will therefore affect large-scale modes of the 21 cm signal along the line of sight. Additionally, the chromatic response of interferometers can cause the foregrounds to contaminate other Fourier modes. Decomposing each wavevector into $\mathbf{k}\equiv(\mathbf{k}_{\perp},k_{\parallel})$, where $k_{\parallel}$ denotes the Fourier modes parallel to the line of sight and $\mathbf{k}_{\perp}$ represents the Fourier modes in the plane perpendicular to the line of sight, then in $_{\perp}-kk_{\parallel}$ space, the region where foreground contamination occurs is referred to as the “foreground wedge” due to its shape Datta et al. (2010); Morales et al. (2012); Parsons et al. (2012); Thyagarajan and others (2015); Liu et al. (2014a, c); Munshi et al. (2025).
@@
In principle, there is more cosmological information encoded in the wedge-obscured modes, because the scaling of the power spectrum and thermal instrumental noise is such that the signal-to-noise ratio is much higher for small wavenumbers. However, it is difficult to completely subtract foregrounds Datta et al. (2010); Liu and Tegmark (2011); Chapman et al. (2012); Ghosh et al. (2015); Ewall-Wice et al. (2021); Mertens et al. (2018); Wang et al. (2024) in the wedge due to uncertainties in modeling both the foreground emission and the chromatic response of the instrument Barry et al. (2016); Patil et al. (2016); Ewall-Wice et al. (2017); Byrne et al. (2019); Mouri Sardarabadi and Koopmans (2019); Neben et al. (2016); Joseph et al. (2018); Ansah-Narh et al. (2018); Zheng and others (2017); Orosz et al. (2019); Joseph et al. (2020); Choudhuri et al. (2021); Kim et al. (2022); Ewall-Wice et al. (2016); Kern et al. (2019); Ung et al. (2020); Josaitis et al. (2022); Rath et al. (2024); Chen et al. (2025b). Therefore, these modes are often discarded, with many analyses of 21 cm data focusing on evading foreground-contaminated modes Vedantham et al. (2012); Trott et al. (2012); Hazelton et al. (2013); Thyagarajan and others (2013); Liu et al. (2014b). This effectively renders a large number of cosmological modes inaccessible to current experiments, significantly reducing their sensitivity. It also limits the possibility of correlating the 21 cm signal with complementary tracers of the matter field from high-redshift galaxy surveys or other line intensity mapping experiments La Plante et al. (2023); Gagnon-Hartman et al. (2025); Fronenberg and Liu (2024); Chen et al. (2025a).
@@
The rest of this paper is organized as follows. In Section II, we provide a conceptual overview of how we reconstruct missing modes and describe the various assumptions we make about what density information is available. In Section III, we briefly review the effective 21 cm bias expansion of Refs. McQuinn and D’Aloisio (2018); Qin et al. (2022) and describe our methods for reconstructing missing density modes. In Section V, we present the results of applying the bias expansion to different combinations of density information, and also demonstrate the complementarity of measurements from galaxy surveys. We conclude in Section VI. Throughout this article, we use cosmological parameters from Planck 2018 Aghanim and others (2020).
@@ II Conceptual Overview
Devising an optimal strategy for extracting density information from 21 cm fluctuations would be worth its own dedicated study, and therefore we choose to focus primarily on steps 3 and 4 in this work. However, we note that Ref. Chen et al. (2025c) has shown that step 2 is possible by adopting a gradient-based sampler to simultaneously infer the bias parameters and the underlying density field from the observed 21 cm modes outside the foreground wedge. Specifically, deep in the perturbative regime, the 21 cm field and density field are primarily related through the linear bias term which is local Qin et al. (2022, 2025). This indicates that the density modes that can be recovered correspond to the same modes where one has uncontaminated 21 cm information. This is corroborated by the findings of Ref. Chen et al. (2025c), which showed explicitly that density modes outside the foreground wedge can be reliably inferred with up to $90\%$ accuracy on large scales (depending on the region of Fourier space) assuming a $\sim 3\sigma$ detection of those same individual 21 cm modes. As it turns out, the same approach of Ref. Chen et al. (2025c) can also be used to reconstruct some of the large-scale density modes in the foreground wedge because their forward model includes nonlinear mode-coupling. These results and future improvements to step 2 can interface with the framework presented here in a straightforward way.
@@
  • the Fourier modes $\mathbf{k}$ outside of the foreground wedge, which are those with Liu et al. (2014b)
@@
  • the modes $k_{\parallel}>k_{\mathrm{cut}}$, corresponding to a scenario where densely spaced or specifically arranged interferometer array configurations make it possible to use foreground mitigation strategies that may be able to recover wedge modes at high $k_{\perp}$ Murray and Trott (2018); MacKay et al. (2025). We also include this case in our discussion because it enables a more direct comparison with results from the literature Darwish et al. (2021). We will refer to this as the “flat” case.
@@ IV Setup and mock fields
To evaluate the expression in Eq. (18), we first filter the nonlinear small scales from the tracer field. As above, we denote the smoothed tracer field by $\delta_{\mathrm{NL}}^{W}$. We calculate the wavenumber cutoff $k_{\mathrm{NL}}$ for this filter by smoothing $\delta_{\mathrm{NL}}$ with a Gaussian window function on increasingly large length scales until the largest overdensity is less than 0.8; the value of $k_{\mathrm{NL}}$ is not particularly sensitive to the choice of value that we use for largest overdensity or the choice of window function Qin et al. (2022). For the 21cmFAST realization used in this work, we find that the nonlinear scale is approximately $k_{\mathrm{NL}}=0.4$ cMpc-1 at z=8.5. The third column of Table 1 shows the number of uncontaminated modes that are in the perturbative regime, i.e. with $k<k_{\mathrm{NL}}$. We note that we have chosen the resolution of our 21cmFAST simulation to sample up to a scale that is twice as large as the nonlinear scale (i.e., $k_{\mathrm{max}}\sim 0.8\ \mathrm{cMpc^{-1}}$ for our simulation). To evaluate the total power spectra in Eq. (23), we use the CAMB Boltzmann code Lewis and Challinor (2011).
@@ V.1 Individual terms
Figure 5: The 21 cm intensity field resulting from different assumptions about the density field substituted into Eq. (12) and smoothed over $k>0.1$ cMpc-1. In order to facilitate comparison between the different input density fields, we consistently use the bias coefficients fit using the full density field and normalize each panel arbitrarily.
Figure 6: Metrics of success for reconstructing the 21 cm signal. The first row shows the cross-correlation with the true 21 cm signal, $r_{\mathrm{21}}$, and the second row shows the ratio of the power spectra of the two fields. For the last three columns, we assume $\delta_{\mathrm{m,obs}}$ are the density modes lying outside of a wedge-shaped foreground region. In order to facilitate comparison of the different input density fields, we use the bias coefficients fit using the full density field across all columns.
@@ V.2 Adding density information from galaxy surveys
In Fig. 5, we show the 21 cm intensity field calculated assuming wedge-shaped foregrounds using different assumptions about the underlying density field and smoothed over $k>0.1$ cMpc-1 (which is on the same order as $k_{\mathrm{cut}}$). In order to ensure that differences between each panel are not merely due to changes in the best-fit bias parameters, in each panel we use the same bias parameters as determined from the full density field. Each panel also has an arbitrary normalization in order to facilitate visual comparison of features. The first panel shows the “true” 21cmFAST field that we are attempting to reconstruct. The second panel shows the result of substituting the full density field, $\delta_{\mathrm{m,full}}$, into Eq. (12). This is the best recovery of the 21 cm signal that one could ever obtain using the bias expansion, and there is indeed a very good resemblance to the true signal. The third and fourth columns serve as a comparison of the total reconstructed 21 cm field in the wedge before and after performing density reconstruction in the wedge and adding this information to observed modes. The phases of the 21 cm modes are not as well reconstructed upon performing density reconstruction in the wedge, hence we do not see the same features as in the first and second panels. The last column of Fig. 6 shows the result of substituting $\delta_{\mathrm{m,obs}}\oplus\delta_{\mathrm{m,Roman}}$ into Eq. (12). In this case, the reconstruction is good enough that we again see prominent 21-cm intensity features in the same regions as in the first panel.
In Fig. 6, we quantify the level of similarity between the true and reconstructed signals, and show $r_{21}$ and $P_{\mathrm{total}}/P_{21}$ for the same sets of assumptions as in Fig. 5. In the first column, we see that the cross-correlation is about 0.9 for wavenumbers within the perturbative regime. The amplitude of the power spectrum is also recovered with high accuracy and is within a factor of 1.5 of the true 21 cm power spectrum.
@@
Since the middle two columns show results for the total reconstructed 21 cm field in the wedge with and without performing density reconstruction in the wedge, they are quite similar to the final columns of Figs. 2 and Figs. 3, but using a slightly different set of bias parameters. As discussed in Section V.1, the cross-correlation is slightly better without first performing density reconstruction. In other words, even if we assume that the matter density in the foreground wedge is zero, which is clearly incorrect, the mode-coupling intrinsic to the EFT terms is enough to yield a cross-correlation at the level of $r\sim 0.3$. This behavior can be understood as arising from the relative contribution of the linear terms with and without in-wedge density reconstruction, as described in Section V.1: the slight degradation of the cross-correlation is due to the fact that the linear terms in the bias expansion are less correlated with the true 21 cm fluctuations than the nonlinear terms. Additionally, without first performing density reconstruction, the reconstructed in-wedge 21 cm field underestimates the power spectrum by a factor of 8, due to the fact that we are missing contributions to the 21 cm fluctuations from many of the underlying density modes, whereas with density reconstruction the reconstructed 21 cm field overestimates the power spectrum by a factor of 90 (this is larger than what was reported in the previous section due to the different bias coefficients).
@@
In the last column, for modes that are within the foreground wedge and also not covered by the galaxy survey, the recovered cross-correlation is at the level of 0.5, and the power spectrum is within a factor of a few of the true amplitude. This is a clear improvement over using either $\delta_{\mathrm{m,obs}}$ or $\delta_{\mathrm{m,obs}}\oplus\delta_{\mathrm{m,recon}}$. This highlights the complementarity of high redshift galaxy surveys to 21cm experiments, and motivates their joint use in recovering the underlying density field as well as reconstruction of 21 cm modes obscured by foregrounds.
@@ VI Conclusion
Assuming that we have a method of obtaining density information outside of the foreground wedge Chen et al. (2025c), we find that simply applying our bias expansion to these modes already yields $\sim 30\%$ cross-correlations with the true 21 cm signal within the foreground region, due to non-linear couplings between small- and large-scale modes. Applying the density reconstruction method of Ref. Darwish et al. (2021) to obtain additional density information within the foreground wedge improves the recovery of linear contributions to the 21 cm signal within the foreground region. For the wedge-shaped foreground scenario, density reconstruction does not improve the overall cross-correlation. Density reconstruction applied to the “flat” and “isotropic” scenarios yields results that are much improved compared to the density-reconstructed “wedge” case, giving cross-correlations as high as $70\%$ and power spectra that are within a factor of a few of the true value—this motivates further study of independent methods for extracting small-scale density information.
@@
In summary, we have demonstrated that bias expansions combined with EFT-based treatments of nonlinear mode-coupling provide a useful starting point for recovering 21 cm information lost to foregrounds. Their utility is enhanced when combined with reliable extractions of the underlying density field within the foreground region, e.g. from galaxy surveys. There are a number of potential directions for improvement on the results presented here. For example, the density estimator of Ref. Darwish et al. (2021) that we apply throughout is derived using an unrenormalized bias expansion; it would be instructive to rederive this estimator using renormalized operators and including all of the appropriate counterterms. Moreover, an assumption that we make throughout this work is that one can extract density information outside of the 21 cm foreground wedge from 21 cm measurements due to the strength of the linear bias term. However, this assumption merits a detailed study in its own right and new techniques to extract this density information, such as those presented in Ref. Chen et al. (2025c), could be combined with the techniques studied here in a straightforward way. We leave these as directions for future study.